What is the correct formula for the isoquant associated with K=10 and L=10


1
Expert's answer
2021-02-02T09:28:35-0500

Solution:

Cobb Douglas production function: Q = 10K0.3L0.6


The general formula of an isoquant in this production function is: "K=\frac{Q}{L}"

First derive Q: Where K = 10, and L = 10

Q = "(10) (10^{0.3} ) \times (10^{0.6} )"

 

= "19.95 \times 3.98 = 79.43"

Q = 79.43

Therefore, the correct isoquant formula:

"K=\frac{Q}{L}"


"K=\frac{79.43}{L^{0.6} }"